MANCED CELL BLOY L.. Rowan University-CHEM09240 - Fall 19 - BANDEGE Chec + Ch Caroline Alemany BANDEGI > Activities and Due D all parts
Answer #1

со мед, оң, уа, +50 — це,04м-той е седем, то е целом. В., roыг * first reaction & #Ney of Not too - NCANA FONT to O. 2 19 м भҚ, = самоң, оң, ми. ) [он] 1ң, оң Сң, Nң. ] 1x 10-*- 2.x О. 29 м-х for weak acid det 0.219M-2 ~0.2299 Пх10-8 м = х2 0-29 м ПхI Part -йон, чи, ти, о 9 недеяч, оғ 1.2.3 хоч (423xto ел. - ко. - 149ся, ст. 9) (o=) E e Car Cry Note) Кь, = х2 again ignowr9-928 грқ, 2 (9 P% - 19- 9928 Рк» = чо+2 Кь, – L0-Ч-o+2 Қ, з 3-4 х (o-se From Equation 3-4ко би= 0 2* 9, 1:23 х х- - 8-ч хtorLor = 2, +22 z 1.23 x 10-44 + 1.01 X10404 = 2:24 X10-4 M pon = -log[2.24x604) = -log 2.24 +E logole = -0.3511 tu . 3.64 pH +pa = 10.35 HM CH₂ Cola No₂ = 0.2188 M q Cola CH₂ Nota = 1.23 810-4 - X2 - 1.23 X164 - 101210-4ry 2 0-2 2 x 104 M. = 1:01 810-